OpenStudy (anonymous):

Please help me!! (I WILL MEDAL/FAN) For f(x) = 0.01(2)^x, find the average rate of change from x = 3 to x = 8.

Vocaloid (vocaloid):

average rate of change from x = 8 to x = 3 can be expressed with the formula averate rate of change = [f(8)-f(3)]/(8-3)

OpenStudy (anonymous):

f(8)-f(3)/5 ??? @Vocaloid sorry was talking to my mum

Vocaloid (vocaloid):

almost done, now we need to find f(8)-f(5) what is f(8)?

OpenStudy (anonymous):

umm..... 8??*_* (Confused)

Vocaloid (vocaloid):

f(x) = 0.01(2)^x, right? f(8) = ?

Vocaloid (vocaloid):

f(8) means find f(x) when x = 8

OpenStudy (anonymous):

f(8)=.01(2)^8

Vocaloid (vocaloid):

right, can you type that out onto a calculator please and tell me what you get?

OpenStudy (anonymous):

2.56

Vocaloid (vocaloid):

right! now, can you find f(3) using the same pattern?

OpenStudy (anonymous):

.08

Vocaloid (vocaloid):

right! f(8) = 2.56 f(3) = 0.08 so what is f(8)-f(3)?

OpenStudy (anonymous):

2.48

Vocaloid (vocaloid):

right! final step: [f(8)-f(3)]/5 = ?

OpenStudy (anonymous):

.496

Vocaloid (vocaloid):

yup, we're done!

OpenStudy (anonymous):

thanks youuuu!!! :)